1. For the Cobb-Douglas production function Y=AKaL1-a, the marginal product of capital is:

a. (1-a)(Y/L)

b. a(Y/K)

c. a

d. aY

2. The principal economic cost of growth is:

a. higher interest rates

b. consumption sacrificed for capital formation

c. higher unemployment rates

d. higher inflation rates

Countries with a high level of real GDP per person are likely to have:

a. low levels of commercial energy use

b. superior literacy levels for females compared to males

c. females with a lower level of life expectancy at birth than males

d. achieved high standards of sanitation
Answer #1

Answer:

(1) Given, the Cobb-Douglas production function Y = AKL1-a

Therefore, the marginal product of capital is дү МРк - әк-әк а О АКL1-а = а.АК-11-а = а.АК“[1-ак-1 = a

Thus option b is the correct answer.

(2) The principal economic cost of growth is higher inflation rates because to attain economic growth a country need to increase real GDPor real income which means Aggregate Demand (AD) increases faster than Aggregate Supply (AS) which again leads to higher inflation as firms start to increase the prices.

Thus, option d is the correct answer.

(3) Countries with a high level of real GDP per person are likely to have achieved high standards of sanitation because the rich people can afford better sanitation and drinking water and health facilities.

Thus option d is the correct answer.